OpenStudy (anonymous):

A number cube with the numbers 1, 2, 3, 4, 5, and 6 is rolled 1000 times. About how many times would it be expected that a number greater than 1 is rolled? A. 167 B. 333 C. 667 D. 833

OpenStudy (anonymous):

p(1)=1/6 p(not 1)=1-(1/6)=5/6 required no. oftimes=1000*(5/6)=5000/6=2500/3=833 (round off)